OpenStudy (anonymous):

find the zeros -4 and 4 and vertex at (0, -16)

OpenStudy (raden):

you mean find the quadratic function which has zeroes .... right ?

OpenStudy (anonymous):

yes the quadratic function with zeros -4 and 4 and vertex at 0, -16

OpenStudy (raden):

use the formula : y = a(x-a)(x-b) with a,b are the zeroes (here known a=-4 and b=4) put a,b into equation above : y = a(x-a)(x-b) y = a(x-(-4))(x-4) y = a(x+4)(x-4) now, subtitute the vertex point (0,-16), giving us y = a(x+4)(x-4) -16 = a(0+4)(0-4) -16 = a(4)(-4) -16 = -16a thus a = -16/-16 = 1 therefore, it is y = a(x-a)(x-b) = 1(x+4)(x-4) = (x+4)(x-4) = x^2 - 16

OpenStudy (anonymous):

i am trying to answer this problem and the options are y=x squared -16, y= -x squared -16, y= x squared +16, or y= -x squared +16

OpenStudy (anonymous):

so the answer is a i see know how you came up with that thank you very much.

OpenStudy (raden):

yup welcome :)

OpenStudy (anonymous):

how do you find the domain of y= 1/x-5 +7